OpenStudy (anonymous):

12. A triangle has sides of lengths 6, 8, and 10. Is it a right triangle? Explain. (2 points)

OpenStudy (anonymous):

True because 6^2 + 8^2 = 10^2

OpenStudy (anonymous):

Plug those values into the pythagorean theorem and see if they're equal. \[6^{2}+8^{2}=10^{2}\] 36 + 64 = 100 100 = 100 So it is a right triangle.

OpenStudy (anonymous):

@seiga is entirely correct. However, if you notice that you can simplify it down to a 3 - 4 - 5 triangle, that is another method of finding these without squaring things.

OpenStudy (anonymous):

İt is not definite that this triangle has a 90 degree angle value. So you cant use those expressions. There is a method.

OpenStudy (anonymous):

@yunus a 3-4-5 triangle is always a right triangle.

OpenStudy (anonymous):

exactly...

OpenStudy (anonymous):

Here i tried to explain.

OpenStudy (anonymous):

@saze i know it is a triangle. everyone knows this. but to proof this you cant directly use the pisagor equation. you should first proof that the valuse of an angle is 90 degree and to do this you can use cosinus theorem.

OpenStudy (anonymous):

But a short way to cheek whether its a triangle or not is to use the triangle inequality which i tried to explain above.